OpenStudy (anonymous):

What is the best way to classify each equation? Column A Column B 1. 2. 3. 4. A. identity B. contradiction C. neither

OpenStudy (anonymous):

@AndrewTheCookie

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

I'll give medals please help

OpenStudy (anonymous):

cookie you come to helkp xD

OpenStudy (anonymous):

help?

OpenStudy (andrewthecookie):

I spy cx

OpenStudy (anonymous):

explane how you got the answer's please

OpenStudy (andrewthecookie):

Im not answering right now

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@Werewolfprincess

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@Daniel_Chernioglo

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

You would just try and solve for x. If you get x equal to one answer ( x=4 or something), then it is "neither" If you end up with true equation ( 0=0 or something), then it's "identity." Lastly if you end up with false equation ( 1=0 ), then it's "contradiction."

OpenStudy (anonymous):

ya i need help solving some of them i don't need help with number 3 though

OpenStudy (anonymous):

i need help with solving number 1 can you help me with that one please?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

@tejasvir

geerky42 (geerky42):

What did you get for number 1? What have you tried so far?

OpenStudy (anonymous):

umm i did 22 - 3x + 7x = 4(x+5) 22 - 3x + 7x = 4x+20

OpenStudy (anonymous):

and that's pretty much it

OpenStudy (anonymous):

how would i solve 22- 3x + 7x

OpenStudy (daniel_chernioglo):

i think you would + 3x to 4x which would= 7x so...22+7x=7x+20

OpenStudy (daniel_chernioglo):

then maybe subtract 7x from 7x→22=x+20

OpenStudy (daniel_chernioglo):

then - 20 from 22 2=x

OpenStudy (anonymous):

lost

OpenStudy (anonymous):

so that one would be neither?

geerky42 (geerky42):

Or try to combine like terms. \(-3x+7x \) is \(4x\) So you have \(22+4x = 4x+20\) Then subtract both sides by 4x. \(22+4x \color{red}{-4x}= 4x\color{red}{-4x}+20\) \(22=20\)

geerky42 (geerky42):

So that's contradiction.

OpenStudy (anonymous):

how woulkd it be contradiction if its equal contradiction dosen't mean equal

OpenStudy (anonymous):

identity means equal

geerky42 (geerky42):

No, identity means equation is true for any values of x. \(2x = 2x\) is identity, because no matter what value you plug in for x, equation will always be true. Contradiction means equation is false for any values of x. Neither means equation is true for certain values of x.

geerky42 (geerky42):

Tell me, is \(22=20\) true or false?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

false

geerky42 (geerky42):

Right, so #1 is contradiction.

OpenStudy (anonymous):

how?

geerky42 (geerky42):

22=20 is false.

OpenStudy (anonymous):

where did you get 22 = 20 from?

geerky42 (geerky42):

Did you see my work above here?

OpenStudy (anonymous):

i didn't understand it

geerky42 (geerky42):

Which part?

OpenStudy (anonymous):

you put

OpenStudy (anonymous):

-3x + 7x is 4x

geerky42 (geerky42):

Yeah, another way to see is that \(-3x+7x = x(-3+7)\), right? And we know that \(-3+7\) is equal to \(4\), so \(-3x+7x = x(-3+7) = 4x\)

OpenStudy (anonymous):

okay

geerky42 (geerky42):

What don't you understand?

OpenStudy (anonymous):

okay now i get that back to the next thing you did

OpenStudy (anonymous):

was that one like a example?

OpenStudy (anonymous):

the one you just did

geerky42 (geerky42):

sorta, we were focus only on \(-3x+7x\) part. So seeing equation as whole, we go from \(22 - 3x + 7x = 4x+20\) to \(22 +4x = 4x+20\)

geerky42 (geerky42):

Now we have 4x in both sides, right? So we just cancel them out, and are left with \(22=20\)

OpenStudy (anonymous):

k thanx i masted on the 5th one all i needed help was on the first one and your right its a contradiction xD BYEEE!

OpenStudy (anonymous):

MEDAL FOR YOU!

geerky42 (geerky42):

Ok glad I helped